# Gravitational Waves in a Perturbative Bianchi I Background

### v1

@@ III Luminosity Distance
Beginning with the only nonzero Christoffel symbols,
$\Gamma^{x}_{tx}=\frac{a^{\prime}}{a},~{}\Gamma^{y}_{ty}=\frac{b^{\prime}}{b},~{}~{}\Gamma^{z}_{tz}=\frac{c^{\prime}}{c},$
$\Gamma^{t}_{xx}=aa^{\prime},~{}~{}\Gamma^{t}_{yy}=bb^{\prime},~{}~{}\Gamma^{t}_{zz}=cc^{\prime},$ (32)
and applying the geodesic equation,
$\Gamma^{\lambda}_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}+\ddot{x}^{\lambda}=0,$ (33)
we get
$aa^{\prime}\dot{x}^{2}+bb^{\prime}\dot{y}^{2}+cc^{\prime}\dot{z}^{2}+\ddot{t}$ $=0$
$2\frac{a^{\prime}}{a}\dot{x}\dot{t}+\ddot{x}$ $=0$
$2\frac{b^{\prime}}{b}\dot{y}\dot{t}+\ddot{y}$ $=0$
$2\frac{c^{\prime}}{c}\dot{z}\dot{t}+\ddot{z}$ $=0,$ (34)
where $\dot{}$ indicates $\frac{d}{dp}$, the first derivative with respect to the affine parameter $p$, and indicates the first derivative with respect to time, $\frac{d}{dt}$. These are related by $\frac{d}{dp}=\dot{t}\frac{d}{dt}$.
The 2nd through 4th equations of Eq. (III) are solved by
$\dot{x}$ $=\frac{A}{a^{2}}$
$\dot{y}$ $=\frac{B}{b^{2}}$
$\dot{z}$ $=\frac{C}{c^{2}}$ (35)
respectively, where $A$, $B$, and $C$ are constants. To first order in $\beta$ and $\gamma$, the second and third lines of Eq. (III) are
$2\frac{a^{\prime}}{a}\dot{y}\dot{t}+\ddot{y}$ $\approx 0$
$2\frac{a^{\prime}}{a}\dot{z}\dot{t}+\ddot{z}$ $\approx 0.$ (36)
Therefore, Eq. (III) shows that $y$ and $z$ satisfy the same equation that $x$ satisfies, so
$\frac{A}{a^{2}}=\frac{B}{b^{2}}=\frac{C}{c^{2}},$ (37)
and thus $B$ and $C$ can be written in terms of $A$ in this way:
$B$ $=A(\frac{b}{a})^{2}$
$C$ $=A(\frac{c}{a})^{2}.$ (38)
Redshift in the Bianchi I metric is approximately
$\mathrm{z}\approx\frac{W(a(t_{0}))}{W(a(t_{i}))}-1,$ (39)
where $W$ is the anisotropic stand-in for the FLRW $a(t)$, defined as Schucker:2014wca
$W(a)=(\frac{A^{2}}{a^{2}}+\frac{B^{2}}{b(a)^{2}}+\frac{C^{2}}{c(a)^{2}})^{-1/2}.$ (40)
There has been some confusion in the past concerning the luminosity distance for the Bianchi I metric, and Schucker:2014wca has clarified this and rederived the luminosity distance for the Bianchi I metric, $d_{BI}$:
$\frac{1}{d_{BI}^{2}}=\frac{4\pi l}{L}=\frac{W(a(t_{i}))^{5}}{a(t_{i})b(t_{i})c(t_{i})W(a(t_{0}))^{4}}\left(\frac{1}{A^{2}I_{y}I_{z}+B^{2}I_{x}I_{z}+C^{2}I_{x}I_{y}}\right),$ (41)
where $l$ is the apparent luminosity (or flux) from an emitting object, $L$ is the absolute luminosity, and $I_{x},I_{y},I_{z}$ are defined as
$I_{x}=-\int_{t_{i}}^{t_{0}}\!\frac{W(a(t))^{3}}{b(a(t))^{2}c(a(t))^{2}}\,\mathrm{d}t,$ (42)
$I_{y}=-\int_{t_{i}}^{t_{0}}\!\frac{W(a(t))^{3}}{a(t)^{2}c(a(t))^{2}}\,\mathrm{d}t,$ (43)
$I_{x}=-\int_{t_{i}}^{t_{0}}\!\frac{W(a(t))^{3}}{a(t)^{2}b(a(t))^{2}}\,\mathrm{d}t.$ (44)
Now, substituting our perturbative forms for $b(a)$ and $c(a)$ from Eq. (3) and expanding to first order in $\beta$ and $\gamma$, we see that the perturbative Bianchi I luminosity distance $d_{PBI}$ is given by
$\frac{1}{d_{PBI}}\approx\frac{(1+\beta+\gamma)}{a(t_{0})\tilde{r}}.$ (45)
This implies
$\frac{1}{a(t_{0})\tilde{r}}\approx\frac{1}{(1+\beta+\gamma)d_{PBI}}.$ (46)
@@
$h_{PBI,+}=\frac{4}{(1+\beta+\gamma)d_{PBI}}(1+\mathrm{z})^{2/3}(GM_{c})^{5/24}(\frac{5}{256}\frac{1}{\tau_{obs}})^{1/4}\cos{(-2(5GM_{c})^{-5/8}\tau_{obs}^{5/8}+\Phi_{0})}\frac{1+\cos^{2}\iota}{2},$ (47)
$h_{PBI,\times}=\frac{4}{(1+\beta+\gamma)d_{PBI}}(1+\mathrm{z})^{2/3}(GM_{c})^{5/24}(\frac{5}{256}\frac{1}{\tau_{obs}})^{1/4}\sin{(-2(5GM_{c})^{-5/8}\tau_{obs}^{5/8}+\Phi_{0})}\cos\iota.$ (48)
If one assumes an FLRW background when fitting a GW, and if our universe is more accurately described with a PBI background, the inference to the fitted luminosity distance would be off by a factor of
$\epsilon=1+\beta+\gamma$ (49)
according to Eqs. (30) and (31), which in turn would bias the inference of the Hubble parameter. This difference in the waveform would manifest as a difference in the size of the envelope surrounding the oscillatory wave (see Fig. 1).
Figure 1: Example gravitational waveform with the envelope indicated.
@@ IV Discussion and Results
$\beta=-(0.011\pm 0.011),$ (50)
$\gamma=0.018\pm 0.014.$ (51)
These constraints imply that the factor difference is
$\epsilon=1.007\pm 0.025.$ (52)
This result means that inferences of luminosity distance could be incorrect by a factor of up to 1.032, which would manifest as a decrease in amplitude of the wave by a factor of 1/1.032.
@@ V Conclusion
In this work, we assumed a perturbative Bianchi I background which was conducive to direct comparison with the FLRW GW solution for a binary system in circular orbit to first order in $v/c$ which includes back-reaction with respect to an FLRW background (”local wave” approximation). We formulated the luminosity distance in this framework and found that inferring the luminosity distance from a GW event may result in error if the wrong background spacetime is assumed for the model fit, as much as up to 3.2%.
Assuming an incorrect background spacetime through which GWs propagate to our observatories has a very small effect on the waveform in general, but the cumulative effect of inferring a slightly biased luminosity distance for several observations may add up to an bad inference of Hubble parameter which is significantly erroneous.

### v2

@@ III Luminosity Distance
The luminosity distance in this metric mimics the same form as in the FLRW metric:
$\frac{\xi}{d_{PBI}(\mathrm{z})}\approx\frac{1}{a(\eta_{0})\tilde{r}(1+\mathrm{z})}.$ (32)
Notice that $\tilde{r}$ plays the role of $r$, and there is a conversion factor $\xi$ which accounts for the fact that comoving distance and luminosity distance are altered relative to their FLRW counterparts. It is defined as
$\xi=\left({\frac{d_{PBI}}{\tilde{r}}}\right)\bigg{/}\left({\frac{d_{FLRW}}{r}}\right).$ (33)
In addition to the perturbative quantities $\beta$ and $\gamma$ that parameterize the anisotropy, $\xi$ is also a function of angle, redshift, and matter density fraction. A universe dominated by matter and a cosmological constant is assumed. The full expression is, using the expression for $d_{PBI}$ from Schucker:2014wca ,
$\xi=$ $\bigg{[}\bigg{(}1+\frac{1}{3(\sqrt{g(\mathrm{z}+1)^{3}+1}-\sqrt{g+1})}\bigg{(}4\left(\frac{\mathrm{z}}{I}-\sqrt{g+1}\right)+\left(\frac{2(\mathrm{z}+1)}{I\sqrt{g(\mathrm{z}+1)^{3}+1}}-3\right)\times$
$\left(\sqrt{g(\mathrm{z}+1)^{3}+1}-\sqrt{g+1}\right)\left((\beta-2\gamma)(\cos^{2}{\theta}-\sin^{2}{\phi}\sin^{2}{\theta})+(\beta+\gamma)\cos{2\phi}\sin^{2}{\theta}\right)\bigg{)}\bigg{)}^{1/2}\times$
$\bigg{(}1-\frac{\sqrt{g(\mathrm{z}+1)^{3}+1}}{3(\sqrt{g(\mathrm{z}+1)^{3}+1}-\sqrt{g+1})\sqrt{g+1}}\bigg{(}(\sqrt{g+1}-\frac{\mathrm{z}}{I})(\beta+\gamma)\;+$
$\frac{\mathrm{z}+1}{I}\left(1-\frac{\sqrt{g+1}}{\sqrt{g(\mathrm{z}+1)^{3}+1}}\right)\left(\beta(1-3\sin^{2}{\phi}\sin^{2}{\theta})+\gamma(1-3\cos^{2}{\theta})\right)\bigg{)}\bigg{)}\bigg{]}^{-1}.$ (34)
$I$ is the elliptic integral
$I=-\int_{0}^{\mathrm{z}}\!\frac{1}{\sqrt{g(\mathrm{z}^{\prime}+1)^{3}+1}}\,\mathrm{d}\mathrm{z}^{\prime},$ (35)
$g$ relates to the the matter density fraction as
$g=\frac{1}{\frac{1}{\Omega_{m}}-1},$ (36)
and $\theta$ and $\phi$ denote angular dependence. $\theta$ is the angle between the source and the $z$ axis, such that a source directly along $+z$ has $\theta=0$ and along $-z$ has $\theta=\pi$. $\phi$ is the azimuthal angle of the source, ranging from 00 to $2\pi$ such that a source in the $x$ direction has $\phi=0$. In the next section we will choose a coordinate system to define these axes on the sky.
@@
$h_{PBI,+}=\xi\frac{GM_{c}(1+\mathrm{z})/c^{2}}{d_{PBI}(\mathrm{z})}\left(\frac{5GM_{c}(1+\mathrm{z})/c^{3}}{t_{coal}-t}\right)^{1/4}\cos{\left(-2\left(\frac{t_{coal}-t}{5GM_{c}(1+\mathrm{z})/c^{3}}\right)^{5/8}+\Phi_{0}\right)}\frac{1+\cos^{2}\iota}{2},$ (37)
$h_{PBI,\times}=\xi\frac{GM_{c}(1+\mathrm{z})/c^{2}}{d_{PBI}(\mathrm{z})}\left(\frac{5GM_{c}(1+\mathrm{z})/c^{3}}{t_{coal}-t}\right)^{1/4}\sin{\left(-2\left(\frac{t_{coal}-t}{5GM_{c}(1+\mathrm{z})/c^{3}}\right)^{5/8}+\Phi_{0}\right)}\cos\iota,$ (38)
Thus, if one assumes an FLRW background when fitting a gravitational wave, and if our universe is more accurately described with a PBI background, the inference to the fitted luminosity distance would be off by a factor of $\xi$ according to Eqs. (30) and (31), which in turn would bias the inference of the Hubble parameter. This difference in the waveform would manifest as a difference in the size of the envelope surrounding the oscillatory wave (see Fig. 1).
Figure 1: Example gravitational waveform with the envelope indicated. We plot Eq. (37), and we use $G=c=1$, $M_{c}=65M_{\odot}$, $t_{coal}=6.41\times 10^{6}M_{\odot}$ ($10^{-6}$ years) after $t=0$, $\mathrm{z}=6.8$, $\Phi_{0}=0$, $\iota=0$, $\xi=0.979$ (its angular-average value, from Eq. (41)), and $d_{PBI}=1.6\times 10^{22}M_{\odot}$ (17 billion light years). For the same luminosity distance as for an FLRW background, the envelope would be affected by a factor of $\xi$. This is true for the case of Eq. (38) as well, which would simply be phase-shifted relative to this plot with these chosen parameters.
@@ IV Discussion and Results
$\beta=-0.69\%,$ (39)
$\gamma=4.29\%.$ (40)
Applying these constraints we obtain an angularly averaged $\xi$ value of
$\bar{\xi}=0.979.$ (41)
This result means that if we live in a PBI universe, inferences of luminosity distance are on average too large, manifesting in gravitational wave amplitudes which are overestimated by $2.1\%$.
Notice that these numerical constraints on anisotropy have at last forced us to choose a coordinate system. The axes have been established as follows:
  • The direction of greatest anisotropy is $+z$.
  • In the plane orthogonal to that axis, the direction of greatest anisotropy is $+x$.
  • $+y$ is now uniquely determined by using a right-handed orthogonal basis.
Perhaps counterintuitively, the values of $\beta$ and $\eta$ do not directly correspond to the anisotropy in the $y$ and $z$ directions. To understand why, recall that the $x$ component of the perturbative Bianchi I metric is not perturbed. Also note that the average value of the Hubble constant must match that of an FLRW universe. Taking these two facts in combination, we see that $a(t)$, the scale factor along the $x$ axis, must take a different value in a PBI universe than in an FLRW one. The anisotropic stand-in for the FLRW $a(t)$ is defined as Schucker:2014wca
$W(a)=\left(\frac{A^{2}}{a^{2}}+\frac{B^{2}}{b(a)^{2}}+\frac{C^{2}}{c(a)^{2}}\right)^{-1/2}.$ (42)
where $A$, $B$, and $C$ are integration constants, and redshift in the Bianchi I metric is approximately
$\mathrm{z}\approx\frac{W(a(t_{0}))}{W(a(t_{i}))}-1,$ (43)
So $H_{F}$ (the isotropic Friedmann Hubble constant) is $\frac{\dot{W}}{W}$ rather than $\frac{\dot{a}}{a}$. It is useful to define a Hubble constant for each direction
$H_{x}$ $=\frac{\dot{a}}{a}$ (44)
$H_{y}$ $=\frac{\dot{b}}{b}$ (45)
$H_{z}$ $=\frac{\dot{c}}{c}$ (46)
with the constraint that
$H_{F}=\frac{H_{x}+H_{y}+H_{z}}{3}.$ (47)
Solving for $\beta$ and $\eta$ in terms of these Hubble parameters reveals the $x$ direction “mixing” that we intuitively argued must exist.
$\beta$ $\propto H_{y}-H_{x}$ (48)
$\gamma$ $\propto H_{z}-H_{x}$ (49)
Fig. (2) shows how these preferred axes related to the Milky Way. If a true anisotropy exists, our universe would be overall “pumpkin shaped” in the orientation shown.
Figure 2: Anisotropically preferred axes plotted against the Milky Way. The white arrow designates the location of the Sun.
The $z$ axis lies almost exactly in the galactic plane and is nearly orthogonal to the Sun-galactic center line. The $x$ axis tips slightly out of the plane Schucker:2014wca . After choosing these axes, thereby concretizing the angles from our $\xi$ expression, we plot the variation of $\xi$ over the sky in Fig. (3).
Figure 3: Angular dependence of $\xi$.
Sources along the $x$ axis would lead to the greatest possible overestimation of luminosity distance, up to 2.8%, and sources along the $z$ axis would have the largest possible underestimation of luminosity distance, up to 5.9% lower than what would be expected in the FLRW case.
@@ V Conclusion
In this work, we assumed a perturbative Bianchi I background which was conducive to direct comparison with the FLRW GW solution for a binary system in circular orbit to first order in $v/c$ which includes back-reaction with respect to an FLRW background (”local wave” approximation). We formulated the luminosity distance in this framework and found that inferring the luminosity distance from a GW event may result in error if the wrong background spacetime is assumed for the model fit, as much as up to 5.9%.
Assuming an incorrect background spacetime through which GWs propagate to our observatories has a very small effect on the waveform in general, but the cumulative effect of inferring a slightly biased luminosity distance for several observations may add up to a bad inference of the Hubble parameter, one that is significantly erroneous. Though the typical luminosity distance uncertainty for current detectors such as LIGO, VIRGO, and KAGRA is 30-60% 061102 ; 1811.12907 ; Chen2018 , future detectors should be able to probe the discrepancy we have derived in this work. Current forecasts for future detectors’ uncertainty in luminosity distance estimation are $\lesssim 10$% for redshifts $\mathrm{z}\lesssim 5$ for the Einstein Telescope and Cosmic Explorer 1912.02622 ; 1907.04833 ; 1206.0331 , and 0.1-10% for the space-based LISA, depending on the source 1511.05581 ; 0906.3752 ; 1601.07112 . Therefore, assuming the wrong background metric will have a measurable effect for future gravitational wave detectors.